OpenStudy (anonymous):

Find an equation of the tangent line to the graph of the function at the point (5, 1). y = 8^(x − 5) can someone please help me

OpenStudy (anonymous):

y=8^(x-5) ln(y)=ln(8^(x-5)) ln(y)=(x-5)ln(8) ln(y)=xln(8)-5ln(8) (1/y)dy/dx=ln(8)-0 dy/dx=y*ln(8) dy/dx=8^(x-5)*ln(8)

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

so is 8^(x-5)ln(8) my equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

its not the answer tho

OpenStudy (anonymous):

o no. i thought u meant something else. thats the equation for ur derivative. u need to use it to find ur slope

OpenStudy (anonymous):

that is the derivative right? you still need the tangent line, so you need the slope. the derivative is \[f'(x)=8^{x-5}\ln(8)\] and \[f'(5)=\ln(8)\]so that is your slope

OpenStudy (anonymous):

yeah lostdove wrote "so that is my equation"

OpenStudy (anonymous):

equation is \[y=\ln(8)(x-5)+1\]

OpenStudy (anonymous):

y=log(8)(x-5)+1?

OpenStudy (anonymous):

is that correct satellite?

OpenStudy (anonymous):

yes its correct

OpenStudy (anonymous):

yes you want the equation of the line. you need the slope and a point, use point-slope formula slope is ln(8) and point is (5,1)

OpenStudy (anonymous):

point slope: if point is \[(x_1,y_1)\] and slope is m then \[y-y_1=m(x-x_1)\]